OpenStudy (bekkah323):

Factor: 5/3x^6+4/3x^5+1/3x^4+1/3x^3

OpenStudy (bekkah323):

\[\frac{ 5 }{ 3 }x^6 + \frac{ 4 }{ 3 }x^5 + \frac{ 1 }{ 3 }x^4 +\frac{ 1 }{ 3 } x^3\]

OpenStudy (luigi0210):

Do you notice anything we can factor?

OpenStudy (bekkah323):

\[\frac{ 1 }{ 3}x^3\]

OpenStudy (luigi0210):

Good, now factor it out and show me what you get.

OpenStudy (bekkah323):

that where i get stuck ( i suck at fractions) i don't know how to divide 5/3 by 1/3

OpenStudy (luigi0210):

Here, maybe doing it this way will help: \[[(5*\frac{ 1 }{ 3 }x^6)+(4*\frac{ 1 }{ 3 }x^5)+\frac{ 1 }{ 3 }x^4+\frac{ 1 }{ 3 }x^3]\]

OpenStudy (bekkah323):

im still confused, sorry! where did the 5/3 and the 4/3 go?

OpenStudy (luigi0210):

I broke them down, doesn't 5*(1/3)=5/3?

OpenStudy (bekkah323):

yes

OpenStudy (luigi0210):

So work backwards to have common factors

OpenStudy (bekkah323):

so 1/3x^3(5x^3+4x^2+x+1)

OpenStudy (luigi0210):

Correct